# Quantum cellular automata and categorical dualities of spin chains

### v1

@@ 6. Examples and Applications
Symmetric QCA. In this section, apply our above analysis to give a better understanding of QCA which commute with a group symmetry, and connect our perspective to the literature. Suppose we have a group $G$ acting unitarily on the local Hilbert space $V:=\mathbbm{C}^{d}$, given by an assignment $g\mapsto U_{g}\in U(d)$. If we let $A$ denote the spin system This extends to the automorphism of each $A_{I}$ by $\alpha_{g}(a):=(U_{g})^{\otimes n}a(U^{*}_{g})^{\otimes n}$, which extends to a homomorphism $G\rightarrow\text{Aut}(A)$. We assume this homomorphism is injective.
@@ Acknowledgments
CJ and KS were supported by NSF DMS-2247202. DJW was supported in part by the Australian Research Council Discovery Early Career Research Award (DE220100625). This paper grew out of discussions that took place during the “QCA group” at the workshop “Fusion categories and tensor networks” at the American Institute of Mathematics in 2022.

### v2

@@ 6. Examples and Applications
Classification of QCA. In this section, apply our above analysis to give a better understanding of QCA which commute with a group symmetry, and connect our perspective to the literature. Suppose we have a group $G$ acting unitarily on the local Hilbert space $V:=\mathbbm{C}^{d}$, given by an assignment $g\mapsto U_{g}\in U(d)$. If we let $A$ denote the spin system This extends to the automorphism of each $A_{I}$ by $\alpha_{g}(a):=(U_{g})^{\otimes n}a(U^{*}_{g})^{\otimes n}$, which extends to a homomorphism $G\rightarrow\text{Aut}(A)$. We assume this homomorphism is injective.
@@
[…]
Let $\alpha$ be a bounded spread automorphism $\alpha\in\text{Aut}(A)$. For any $z\in\mathbbm{Z}$, define $A_{z}\subseteq A$ as the C*-algebra of observables localized in the interval $(-\infty,z]$. Fusion spin chains $A:=A(\mathcal{C},X)$ satisfy the unique trace property, as do all the algebras $A_{z}$.
Now choose $x,y\in\mathbbm{Z}$ so that $y\leq x$, $A_{y}\subseteq\alpha(A_{x})$, which can be accomplished due to the bounded spread condition.
Now we can define the index
$\text{Ind}(\alpha):=\left(\frac{[\alpha(A_{x}):A_{y}]}{[A_{x}:A_{y}]}\right)^{\frac{1}{2}}.$
For ordinary spin chains, this agrees with the GNVW index [JL24, Proposition 4.2]. Our goal is to show that for the group case, if a $G$-duality $\alpha$ duality admits an extension to a QCA $\widetilde{\alpha}$, then $\text{Ind}(\alpha)=\text{Ind}(\widetilde{\alpha})$. This would imply that if $\text{Ind}(\alpha)=1$, $\widetilde{\alpha}$ is a finite depth circuit.
Proposition 6.4.
Let $\widetilde{\alpha}:A\rightarrow A$ be a QCA such that $\widetilde{\alpha}(A^{G})=A^{G}$, and let $\alpha:=\widetilde{\alpha}|_{A^{G}}$ Then $\text{Ind}(\widetilde{\alpha})=\text{Ind}(\alpha)$.
Proof.
Let $x\leq y$ be as in the definition of index. Since the $G$-representation is a strong tensor generator, it follows that $A_{x},A_{y},A^{G}_{x},A^{G}_{y},\widetilde{\alpha}(A_{x}),\alpha(A^{G}_{x})$ are are simple C*-algebras with unique trace. Then we have the inclusions
[…]
Since the indices are actually indices of subfactors, the index is multiplicative [Jon83], hence
$[A_{x}:A^{G}_{x}][A^{G}_{x}:A^{G}_{y}]=[A_{x}:A_{y}][A_{y}:A^{G}_{y}]$
and
$[\widetilde{\alpha}(A_{x}):\alpha(A^{G}_{x})][\alpha(A^{G}_{x}):A^{G}_{y}]=[\widetilde{\alpha}(A_{x}):A_{y}][A_{y}:A^{G}_{y}].$
But from standard results in subfactor theory (for example [JS97, EK98]),
$[A_{x}:A^{G}_{x}]=[A_{y}:A^{G}_{y}]=[\widetilde{\alpha}(A_{x}):\alpha(A^{G}_{x})]=[A_{y}:A^{G}_{y}]=|G|.$
Thus we have $[A^{G}_{x}:A^{G}_{y}]=[A_{x}:A_{y}]$ and $[\alpha(A^{G}_{x}):A^{G}_{y}]=[\widetilde{\alpha}(A_{x}):A_{y}]$. Thus $\text{Ind}_{G}(\alpha)=\text{Ind}(\alpha)$.
Proof of Corollary 1.4. Suppose $\alpha:A^{G}\rightarrow A^{G}$ is a G-duality such that $\text{Ind}(\alpha)=1$ and $\text{DHR}(\alpha)=\text{Id}_{\mathcal{Z}(\text{Rep}(G))}$. Then by Theorem 1.3 there is a lift of $\alpha$ to an sQCA $\widetilde{\alpha}$ which can be chosen by 6.3 to have trivial $H^{2}$ index. Furthermore, by Proposition 6.4, $\text{Ind}(\widetilde{\alpha})=1$, so $\widetilde{\alpha}$ is a finite depth circuit.
@@ Acknowledgments
CJ and KS were supported by NSF DMS-2247202. DJW was supported in part by the Australian Research Council Discovery Early Career Research Award (DE220100625). This paper grew out of discussions that took place during the “QCA group” at the workshop “Fusion categories and tensor networks” at the American Institute of Mathematics in 2022. After completing the first arXiv version of this work we became aware of the related work [MLC24]. We have now included comments on the relationship between our results and those of [MLC24].